Fred is 3 years older than his sister mary. In 7 years she will be six-sevenths of his age. how old are they?
Answer by mananth(16946)

mary's age =x
fred's age will be x+3
..
after 7 years
mary's age = x+7
Fred's age = x+3+7 = x+10
..
x+7= 6/7 * (x+10)
multiply by 7
7(x+7)=(6/7)*7(x+10)
7x+49=6x+60
7x-6x= 60-49
x= 11 years - Mary's age
Fred's age = x+3 = 11+3=14